# Learning Stackable and Skippable LEGO Bricks for Efficient, Reconfigurable, and Variable-Resolution Diffusion Modeling

### v1

@@ Abstract
Diffusion models excel at generating photo-realistic images but come with significant computational costs in both training and sampling. While various techniques address these computational challenges, a less-explored issue is designing an efficient and adaptable network backbone for iterative refinement. Current options like U-Net and Vision Transformer often rely on resource-intensive deep networks and lack the flexibility needed for generating images at variable resolutions or with a smaller network than used in training. This study introduces LEGO bricks, which seamlessly integrate Local-feature Enrichment and Global-content Orchestration. These bricks can be stacked to create a test-time reconfigurable diffusion backbone, allowing selective skipping of bricks to reduce sampling costs and generate higher-resolution images than the training data. LEGO bricks enrich local regions with an MLP and transform them using a Transformer block while maintaining a consistent full-resolution image across all bricks. Experimental results demonstrate that LEGO bricks enhance training efficiency, expedite convergence, and facilitate variable-resolution image generation while maintaining strong generative performance. Moreover, LEGO significantly reduces sampling time compared to other methods, establishing it as a valuable enhancement for diffusion models.
@@ 1 Introduction
Despite these advancements, diffusion models continue to face a significant challenge: their substantial computational requirements, not only during training but also during sampling. In sampling, diffusion models typically demand a large number of functional evaluations (NFE) to simulate the reverse diffusion process. To tackle this challenge, various strategies have been proposed, including methods aimed at expediting sampling through approximated reverse steps (Song et al., 2020; Zhang & Chen, 2022; Lu et al., 2022; Karras et al., 2022), as well as approaches that integrate diffusion with other generative models (Zheng et al., 2022b; Pandey et al., 2022).
@@
Beside the computing-intensive and rigid sampling process, training diffusion models also entails a considerable number of iterations to attain convergence to an acceptable checkpoint. This requirement arises from the model’s need to learn how to predict the mean of clean images conditioned on noisy inputs (Robbins, 1992; Efron, 2011; Luo, 2022; Chung et al., 2022; Zhou et al., 2023), which exhibit varying levels of noise. This mechanism imposes a significant computational burden, particularly when dealing with larger and higher dimensional datasets. To mitigate the training cost, commonly-used approaches include training the diffusion model in a lower-dimensional space and then recovering the high-dimensional data using techniques such as pre-trained auto-encoders (Rombach et al., 2022; Li et al., 2022b) or cascaded super-resolution models (Ho et al., 2022; Saharia et al., 2022). An alternative strategy involves training diffusion models at the patch-level input (Luhman & Luhman, 2022), requiring additional methods to ensure generated samples maintain a coherent global structure. These methods may include incorporating full-size images periodically (Wang et al., 2023), using encoded or downsampled semantic features (Ding et al., 2023; Arakawa et al., 2023), or employing masked transformer prediction (Zheng et al., 2023; Gao et al., 2023).
Figure 1: Left Panel: Comparison of convergence measured with FID (with classifier-free guidance) vs. training time. Right Panel: Comparison of the computation cost measured with FID (with classifier-free guidance) vs. training FLOPs. The experiments are conducted with eight NVIDIA A100 GPUs on ImageNet ($256\times 256$), with a batch size of 256.
@@
In summary, the stackable and skippable LEGO bricks possess several noteworthy characteristics in managing computational costs in both training and sampling, offering a flexible and test-time reconfigurable backbone, and facilitating variable-resolution generation. Our experimental evaluations, as shown in Figures 1 and 2, clearly demonstrate that LEGO strikes a good balance between efficiency and performance across challenging image benchmarks. LEGO significantly enhances training efficiency, as evidenced by reduced FLOPs, faster convergence, and shorter training times, all while maintaining a robust generative performance. These advantages extend seamlessly into the sampling phase, where LEGO achieves a noteworthy 60% reduction in sampling time compared to DiT (Peebles & Xie, 2022), while keeping the same NFE. Additionally, LEGO has the capability to generate images at much higher resolutions ($e.g.$, $2048\times 600$) than training images ($e.g.$, $256\times 256$).
2 Related work
Diffusion models have garnered increasing attention from researchers seeking to enhance them in various ways. These efforts include combining variational loss with weighted denoising loss (Dhariwal & Nichol, 2021; Kingma et al., 2021; Karras et al., 2022), accelerating sampling speed using ODE solvers (Kong & Ping, 2021; San-Roman et al., 2021; Song et al., 2020; Zhang & Chen, 2022; Lu et al., 2022), and employing auxiliary models to improve training (Rombach et al., 2022; Li et al., 2022b; Zheng et al., 2022b; Pandey et al., 2022). Additionally, the iterative refinement property of diffusion models has inspired research in various domains, such as stable GAN training (Xiao et al., 2022; Wang et al., 2022), large-scale text-to-image models (Kim & Ye, 2021; Nichol et al., 2021; Gu et al., 2022; Saharia et al., 2022), and robust super-resolution and editing models (Meng et al., 2021; Choi et al., 2021; Li et al., 2022a), among others.
@@ 3 Local-feature Enrichment and Global-content Orchestration
While the downsampling and upsampling convolutional layers in U-Net allow it to effectively capture local structures, achieving a good understanding of global structures often requires a deep architecture, resulting in high computational complexity during both training and generation. Instead of depending on a large number of convolutional layers, a classical approach in signal and image processing involves the extraction of local feature descriptors and their application in downstream tasks (Lowe, 2004; Mikolajczyk & Schmid, 2005; Tuytelaars & Mikolajczyk, 2008). For instance, techniques like dictionary learning and sparse coding operate on the principle that each local patch can be expressed as a sparse linear combination of dictionary atoms learned from image patches. This learned patch dictionary can then be used to enrich all overlapping patches for various image-processing tasks, such as denoising and inpainting (Aharon et al., 2006; Mairal et al., 2007; Zhou et al., 2009).
The emergence of ViT in image classification (Dosovitskiy et al., 2021), which integrates extracting feature embeddings from non-overlapping local patches with the global attention mechanism of the Transformer, marks a resurgence of this classical local-feature-based concept. ViT has not only competed effectively with state-of-the-art convolution-based networks in both image classification (Touvron et al., 2022) and generation (Esser et al., 2021), but has also been adapted to replace the U-Net in diffusion models (Yang et al., 2022; Bao et al., 2022; Peebles & Xie, 2022; Gao et al., 2023), achieving state-of-the-art image generation performance. This underscores the notion that the convolutional architecture can be supplanted by the enrichment of local features and their orchestration through global attention.
ViT-based models start by enriching patch-level representations and then utilize a sequence of Transformer blocks to capture dependencies among local patches. They often demand smaller patch sizes to capture fine-grained local details and a large number of attention layers to capture global dependencies. This necessitates lengthy sequence computations and deep networks to achieve optimal results, thereby incurring significant computational costs.
@@ 3.1 Motivations for Constructing LEGO Bricks
Our envisioned LEGO bricks are intended to possess several advantageous properties: 1) Spatial Efficiency in Training: Within the LEGO brick ensemble, the majority of them are dedicated to producing local patches using computation-light MLP mixing and attention modules. This design choice leads to a significant reduction in computational Floating-Point Operations (FLOPs) and substantially shortens the overall training duration. 2) Spatial Efficiency in Sampling: During sampling, the LEGO bricks can be selectively skipped at each time step without a discernible decline in generation performance. Specifically, when $t$ is large, indicating greater uncertainty in the global spatial structure, more patch-level LEGO bricks can be safely skipped. Conversely, when $t$ is small, signifying a more stable global spatial structure, more full-resolution LEGO bricks can be bypassed. 3) Versatility: LEGO bricks showcase remarkable versatility, accommodating both end-to-end training and sequential training from lower to upper bricks, all while enabling generation at resolutions significantly higher than those employed during training. Furthermore, they readily support the integration of existing pre-trained models as LEGO bricks, enhancing the model’s adaptability and ease of use.
@@ 3.2 Technical Preliminary of Diffusion-based Generative Models
Diffusion-based generative models employ a forward diffusion chain to gradually corrupt the data into noise, and an iterative-refinement-based reverse diffusion chain to regenerate the data. The reverse diffusion chain can be optimized by maximizing the evidence lower bound (ELBO) (Blei et al., 2017) of a variational autoencoder (Kingma & Welling, 2014), using a hierarchical prior with $T$ stochastic layers (Ho et al., 2020; Song et al., 2020; 2021; Kingma et al., 2021). Let’s represent the data distribution as $q({\mathbf{x}}_{0})$ and the generative prior as $p({\mathbf{x}}_{T})$. The forward and reverse diffusion processes discretized into $T$ steps can be expressed as:
@@ 3.3.1 LEGO Brick Ensemble and Vertical Spatial Refinement
Recursive ensemble of LEBO bricks: The vanilla denoising diffusion step, as represented in Equation equation 6, is decomposed in LEGO into $K$ consecutive LEGO bricks, stacked from the top to the bottom as follows:
@@
with ${\mathbf{z}}_{t}^{(0)}:=\emptyset$, $\theta:=\{\theta_{k}\}_{1,K}$, and ${\mathbf{z}}^{(k)}_{t}$ denoting a grid of refined patches based on the corresponding patches from the output of the lower LEGO brick at time $t$.
We note that since each LEGO brick starts with a full-resolution image and ends with a refined full-resolution image, we have the flexibility to choose the target brick size for each LEGO brick. In our illustration shown in Figure 3, we follow a progressive-growing-based construction, where the patch size of the stacked LEGO bricks monotonically increases when moving from lower to upper bricks. Alternatively, we can impose a monotonically decreasing constraint, corresponding to progressive refinement. In this work, we explore both progressive growth and refinement as two special examples and present results for both cases. The combinatorial optimization of brick sizes for the LEGO bricks in the stack is a topic that warrants further investigation.
@@
Denote $\epsilon$ as a normal noise used to corrupt the clean image. The clean image patches, denoted as ${\mathbf{x}}_{0,(i,j)}^{(k)}$, are of size $r_{k}\times r_{k}$ and correspond to specific spatial locations $(i,j)$. Additionally, we have refined patches, $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k)}$, produced by the $k^{th}$ LEGO brick at the same locations, and ${\hat{{\mathbf{x}}}}_{0,(i,j)}^{(k-1)}$ from the $(k-1)^{th}$ LEGO brick. When processing a noisy image ${\mathbf{x}}_{t}$ at time $t$, we perform upward propagation through the stacked LEGO bricks to progressively refine the full image. This begins with $\hat{{\mathbf{x}}}_{0}^{(0)}=\emptyset$ and proceeds with refined image patches $\hat{{\mathbf{x}}}_{0}^{(k)}$ for $k=1,\ldots,K$. As each LEGO brick operates on sampled patches, when the number of patches is limited, there may be instances where $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ contain missing values. In such cases, we replace these missing values with the corresponding pixels from ${\mathbf{x}}_{0}$ to ensure the needed $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ is present. Denote $\lambda_{t}^{(k)}$ as time- and brick-dependent weight coefficients, whose settings are described in Appendix B. With the refined image patches $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k)}$, we express the training loss over the $K$ LEGO bricks as
@@ 3.3.2 Network Design within the LEGO Brick
In the design of a LEGO brick, we draw inspiration from the design of the Diffusion Transformer (DiT) proposed in Peebles & Xie (2022). The DiT architecture comprises several key components: Patch Embedding Layer: Initially, the input image is tokenized into a sequence of patch embeddings. DiT Blocks: These are Transformer blocks featuring multi-head attention, zero-initialized adaptive layer norm (adaLN-zero), and MLP layers. Linear MLP with adaLN: Positioned on top of the DiT blocks, this linear MLP, along with adaLN, converts the sequence of embeddings back into an image. The diffusion-specific conditions, such as time embedding and class conditions, are encoded within the embeddings through adaLN in every block.
@@
We categorize the LEGO bricks as ‘patch-bricks’ if their brick size is smaller than the image resolution, and as ‘image-bricks’ otherwise. The patch-bricks have attention spans as short as four, saving memory and computation in both training and generation. While the image-bricks choose longer attention spans and hence do not yield memory or computation savings during training, they can be compressed with fewer DiT blocks compared with the configuration in Peebles & Xie (2022). During generation, LEGO bricks can also be selectively skipped at appropriate time steps without causing clear performance degradation, resulting in substantial savings in generation costs. For simplicity, we maintain a fixed embedding size and vary only the number of DiT blocks in each brick based on $r_{k}$. Each LEGO brick may require a different capacity depending on $r_{k}$. For a comprehensive configuration of the LEGO bricks, please refer to Table 4 in Appendix B.
We explore two spatial refinement settings and stack LEGO bricks accordingly: 1) Progressive Growth (PG): In this setup, we arrange LEGO bricks in a manner where each subsequent brick has a patch size that’s four times as large as the previous one, $i.e.$, $r_{k}=4r_{k-1}$. Consequently, the patch of the current break would aggregate the features of four patches output by the previous brick, facilitating global-content orchestration. 2) Progressive Refinement (PR): By contrast, for the PR configuration, we stack LEGO bricks in reverse order compared to PG, with $r_{k}=r_{k-1}/4$. Here, each LEGO brick is responsible for producing a refined generation based on the content provided by the brick below it. These two settings offer distinct approaches to spatial refinement, and we explore their respective performance in our experiments.
@@ 4 Experiments
We present a series of experiments designed to assess the efficacy and versatility of using stacked LEGO bricks as the backbone for diffusion models. We begin by showcasing results on both small-scale and large-scale datasets to evaluate the image generation performance using the three configurations outlined in Table 4 in Appendix B.
4.1 Experimental settings
Training settings: For $64\times 64\times 3$ images, we directly train in the pixel space. For higher-resolution images, we encode them into a latent space with a down-sampling factor of 8 (Rombach et al., 2022), i.e., preprocessing images of size $256\times 256\times 3$ into $32\times 32\times 4$, and $512\times 512\times 3$ into $64\times 64\times 4$, and then train on that latent data. For our LEGO model backbones, we employ EDM (Karras et al., 2022) preconditioning on pixel-space and iDDPM (Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021) on latent-space data, strictly adhering to the diffusion designs prescribed by these respective preconditioning strategies. For further details, please refer to Appendix B.

### v3

@@ Abstract
Diffusion models excel at generating photo-realistic images but come with significant computational costs in both training and sampling. While various techniques address these computational challenges, a less-explored issue is designing an efficient and adaptable network backbone for iterative refinement. Current options like U-Net and Vision Transformer often rely on resource-intensive deep networks and lack the flexibility needed for generating images at variable resolutions or with a smaller network than used in training. This study introduces LEGO bricks, which seamlessly integrate Local-feature Enrichment and Global-content Orchestration. These bricks can be stacked to create a test-time reconfigurable diffusion backbone, allowing selective skipping of bricks to reduce sampling costs and generate higher-resolution images than the training data. LEGO bricks enrich local regions with an MLP and transform them using a Transformer block while maintaining a consistent full-resolution image across all bricks. Experimental results demonstrate that LEGO bricks enhance training efficiency, expedite convergence, and facilitate variable-resolution image generation while maintaining strong generative performance. Moreover, LEGO significantly reduces sampling time compared to other methods, establishing it as a valuable enhancement for diffusion models. Our code and project page are available at https://jegzheng.github.io/LEGODiffusion.
@@ 1 Introduction
Despite these advancements, diffusion models continue to face a significant challenge: their substantial computational requirements, not only during training but also during sampling. In sampling, diffusion models typically demand a large number of functional evaluations (NFE) to simulate the reverse diffusion process. To tackle this challenge, various strategies have been proposed, including methods aimed at expediting sampling through approximated reverse steps (Song et al., 2020; Zhang & Chen, 2022; Lu et al., 2022; Karras et al., 2022), as well as approaches that integrate diffusion with other generative models (Zheng et al., 2022c; Pandey et al., 2022).
@@
Beside the computing-intensive and rigid sampling process, training diffusion models also entails a considerable number of iterations to attain convergence to an acceptable checkpoint. This requirement arises from the model’s need to learn how to predict the mean of clean images conditioned on noisy inputs, which exhibit varying levels of noise. This concept can be concluded from either the Tweedie’s formula (Robbins, 1992; Efron, 2011), as discussed in Luo (2022) and Chung et al. (2022), or the Bregman divergence (Banerjee et al., 2005), as illustrated in Zhou et al. (2023). This mechanism imposes a significant computational burden, particularly when dealing with larger and higher dimensional datasets. To mitigate the training cost, commonly-used approaches include training the diffusion model in a lower-dimensional space and then recovering the high-dimensional data using techniques such as pre-trained auto-encoders (Rombach et al., 2022; Li et al., 2022b) or cascaded super-resolution models (Ho et al., 2022; Saharia et al., 2022). An alternative strategy involves training diffusion models at the patch-level input (Luhman & Luhman, 2022), requiring additional methods to ensure generated samples maintain a coherent global structure. These methods may include incorporating full-size images periodically (Wang et al., 2023), using encoded or downsampled semantic features (Ding et al., 2023; Arakawa et al., 2023), or employing masked transformer prediction (Zheng et al., 2023; Gao et al., 2023).
@@
Figure 2: With classifier-free guidance, we present the following: Left Panel: A comparison of convergence, measured with FID versus training time. Right Panel: A comparison of the computation cost measured with FID versus training FLOPs. These experiments were conducted using eight NVIDIA A100 GPUs on ImageNet ($256\times 256$), with a batch size of 256.
In summary, the stackable and skippable LEGO bricks possess several noteworthy characteristics in managing computational costs in both training and sampling, offering a flexible and test-time reconfigurable backbone, and facilitating variable-resolution generation. Our experimental evaluations, as shown in Figures 1 and 2, clearly demonstrate that LEGO strikes a good balance between efficiency and performance across challenging image benchmarks. LEGO significantly enhances training efficiency, as evidenced by reduced FLOPs, faster convergence, and shorter training times, all while maintaining a robust generation performance. These advantages extend seamlessly into the sampling phase, where LEGO achieves a noteworthy 60% reduction in sampling time compared to DiT (Peebles & Xie, 2022), while keeping the same NFE. Additionally, LEGO has the capability to generate images at much higher resolutions ($e.g.$, $2048\times 600$) than training images ($e.g.$, $256\times 256$).
@@ 2.1 Motivations for Constructing LEGO Bricks
Our envisioned LEGO bricks are intended to possess several advantageous properties: 1) Spatial Efficiency in Training: Within the ensemble, the majority of LEGO bricks are dedicated to producing local patches using computation-light MLP mixing and attention modules. This design choice leads to a significant reduction in computational Floating-Point Operations (FLOPs) and substantially shortens the overall training duration. 2) Spatial Efficiency in Sampling: During sampling, the LEGO bricks can be selectively skipped at each time step without a discernible decline in generation performance. Specifically, when $t$ is large, indicating greater uncertainty in the global spatial structure, more patch-level LEGO bricks can be safely skipped. Conversely, when $t$ is small, signifying a more stable global spatial structure, more full-resolution LEGO bricks can be bypassed. 3) Versatility: LEGO bricks showcase remarkable versatility, accommodating both end-to-end training and sequential training from lower to upper bricks, all while enabling generation at resolutions significantly higher than those employed during training. Furthermore, they readily support the integration of existing pre-trained models as LEGO bricks, enhancing the model’s adaptability and ease of use.
@@ 2.2 Technical Preliminary of Diffusion-based Generative Models
Diffusion-based generative models employ a forward diffusion chain to gradually corrupt the data into noise, and an iterative-refinement-based reverse diffusion chain to regenerate the data. We use $\alpha_{t}\in[0,1]$, a decreasing sequence, to define the variance schedule $\{\beta_{t}=1-\frac{\alpha_{t}}{\alpha_{t-1}}\}_{t=1}^{T}$ and denote ${\mathbf{x}}_{0}$ as a clean image and ${\mathbf{x}}_{t}$ as a noisy image. The reverse diffusion chain can be optimized by maximizing the evidence lower bound (ELBO) (Blei et al., 2017) of a variational autoencoder (Kingma & Welling, 2014), using a hierarchical prior with $T$ stochastic layers (Ho et al., 2020; Song et al., 2020; 2021; Kingma et al., 2021; Zhou et al., 2023). Let’s represent the data distribution as $q({\mathbf{x}}_{0})$ and the generative prior as $p({\mathbf{x}}_{T})$. The forward and reverse diffusion processes discretized into $T$ steps can be expressed as:
@@ 2.3.1 LEGO Brick Ensemble and Vertical Spatial Refinement
Recursive ensemble of LEGO bricks: The vanilla denoising diffusion step, as in Equation 6, is decomposed into $K$ consecutive LEGO bricks, stacked from the top to the bottom as follows:
@@
with ${\mathbf{z}}_{t}^{(0)}:=\emptyset$, $\theta:=\{\theta_{k}\}_{1,K}$, and ${\mathbf{z}}^{(k)}_{t}$ denoting a grid of refined patches based on the corresponding patches from the output of the lower LEGO brick at time $t$. We note that since each LEGO brick starts with a full-resolution image and ends with a refined full-resolution image, we have the flexibility to choose the target brick size for each LEGO brick. In our illustration shown in Figure 3, we follow a progressive-growing-based construction, where the patch size of the stacked LEGO bricks monotonically increases when moving from lower to upper bricks. Alternatively, we can impose a monotonically decreasing constraint, corresponding to progressive refinement. In this work, we explore both progressive growth and refinement as two special examples, and further present a Unet-inspired structural variant that combines both. The combinatorial optimization of brick sizes for the LEGO bricks in the stack is a topic that warrants further investigation.
@@
Denote $\epsilon$ as a normal noise used to corrupt the clean image. The clean image patches, denoted as ${\mathbf{x}}_{0,(i,j)}^{(k)}$, are of size $r_{k}\times r_{k}$ and correspond to specific spatial locations $(i,j)$. Additionally, we have refined patches, $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k)}$, produced by the $k^{th}$ LEGO brick at the same locations, and ${\hat{{\mathbf{x}}}}_{0,(i,j)}^{(k-1)}$ from the $(k-1)^{th}$ LEGO brick. When processing a noisy image ${\mathbf{x}}_{t}$ at time $t$, we perform upward propagation through the stacked LEGO bricks to progressively refine the full image. This begins with $\hat{{\mathbf{x}}}_{0}^{(0)}=\emptyset$ and proceeds with refined image patches $\hat{{\mathbf{x}}}_{0}^{(k)}$ for $k=1,\ldots,K$. As each LEGO brick operates on sampled patches, when the number of patches is limited, there may be instances where $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ contain missing values. In such cases, we replace these missing values with the corresponding pixels from ${\mathbf{x}}_{0}$ to ensure the needed $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ is present. Denote $\lambda_{t}^{(k)}$ as time- and brick-dependent weight coefficients, whose settings are described in Appendix C. With the refined image patches $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k)}$, we express the training loss over the $K$ LEGO bricks as
@@ 2.3.2 Network Design within the LEGO Brick
In the design of a LEGO brick, we draw inspiration from the design of the Diffusion Transformer (DiT) proposed in Peebles & Xie (2022). The DiT architecture comprises several key components: Patch Embedding Layer: Initially, the input image is tokenized into a sequence of patch embeddings; DiT Blocks: These are Transformer blocks featuring multi-head attention, zero-initialized adaptive layer norm (adaLN-zero), and MLP layers; Linear layer with adaLN: Positioned on top of the DiT blocks, this linear layer, along with adaLN, converts the sequence of embeddings back into an image. The diffusion-specific conditions, such as time embedding and class conditions, are encoded within the embeddings through adaLN in every block.
@@
We categorize the LEGO bricks as ‘patch-bricks’ if their brick size is smaller than the image resolution, and as ‘image-bricks’ otherwise. The patch-bricks have attention spans as short as four, saving memory and computation in both training and generation. While the image-bricks choose longer attention spans and hence do not yield memory or computation savings during training, they can be compressed with fewer DiT blocks compared with the configuration in Peebles & Xie (2022). During generation, LEGO bricks can also be selectively skipped at appropriate time steps without causing clear performance degradation, resulting in substantial savings in generation costs. For simplicity, we maintain a fixed embedding size and vary only the number of DiT blocks in each brick based on $r_{k}$. Each LEGO brick may require a different capacity depending on $r_{k}$. For a comprehensive configuration of the LEGO bricks, please refer to Table 5 in Appendix C.
We explore three spatial refinement settings and stack LEGO bricks accordingly: 1) Progressive Growth (PG): In this setup, we arrange LEGO bricks in a manner where each subsequent brick has a patch size that’s four times as large as the previous one, $i.e.$, $r_{k}=4r_{k-1}$. Consequently, the patch of the current break would aggregate the features of four patches output by the previous brick, facilitating global-content orchestration. 2) Progressive Refinement (PR): By contrast, for the PR configuration, we stack LEGO bricks in reverse order compared to PG, with $r_{k}=r_{k-1}/4$. Here, each LEGO brick is responsible for producing a refined generation based on the content provided by the brick below it. 3) LEGO-U: a hierarchical variant that combines the features of LEGO-PR and LEGO-PG. This model processes image patches starting with larger resolutions, transitioning to smaller ones, and then reverting back to larger resolutions, similar to Unets that operate with multiple downsampling and upsampling stages to have multi-scale representations.
@@ 3 Experiments
We present a series of experiments designed to assess the efficacy and versatility of using stacked LEGO bricks as the backbone for diffusion models. We defer the details on datasets, training settings, and evaluation metrics into Appendix C. We begin by showcasing results on both small-scale and large-scale datasets to evaluate the image generation performance using the three configurations outlined in Table 5 in Appendix C.